# Global Dynamics of McKean-Vlasov SDEs via Stochastic Order

### v1

@@ 5.3 Local attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
(iv). For $k=1$, according to the statement (ii), the open ball $B(\delta_{a_{1}},r_{1})$ is positively invariant under $P^{*}$. By Theorem 2.3 (i)(ii) and Remark 5.1, there is only one invariant measure $\nu_{1}$ in $\overline{B(\delta_{a_{1}},r_{1})}$. Theorem 2.1 (ii) and Theorem 2.3 (i) entail that $[\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{1}},r_{1})$ and hence $\omega(B(\delta_{a_{1}},r_{1}))\subset[\nu_{1},\nu_{2n-1}]$. Then $\nu_{1}\leq_{\mathrm{st}}\omega(B(\delta_{a_{1}},r_{1}))$. Let $\{\nu_{t}^{1,\downarrow}\}_{t\in\mathbb{R}}$ be a decreasing connecting orbit from $\nu_{2}$ to $\nu_{1}$ given in Theorem 2.3 (iii). Proposition 5.9 shows that there exists $t_{1}\in\mathbb{R}$ such that $\nu_{t_{1}}^{1,\downarrow}\geq_{\mathrm{st}}\omega(B(\delta_{a_{1}},r_{1}))$. Then Theorem 5.7 (ii) implies that $\nu_{1}$ attracts $B(\delta_{a_{1}},r_{1})$ under $P^{*}$, which means $\sup_{\mu\in B(\delta_{a_{k}},r_{k})}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$ as $t\to\infty$.
For $k=2,3,\dots,n-1$, let $\{\nu_{t}^{k,\downarrow}\}_{t\in\mathbb{R}}$ be a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$ and $\{\nu_{t}^{k,\uparrow}\}_{t\in\mathbb{R}}$ be an increasing connecting orbit from $\nu_{2k-2}$ to $\nu_{2k-1}$ given in Theorem 2.3 (iii). It follows from the statement (ii) that $B(\delta_{a_{k}},r_{k})$ is positively invariant under $P^{*}$. By Theorem 2.3 (i)(ii) and Remark 5.1, there is only one invariant measure $\nu_{2k-1}$ in $\overline{B(\delta_{a_{k}},r_{k})}$. We also see the order interval $[\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{k}},r_{k})$ by Theorem 2.1 (ii) and Theorem 2.3 (i). By Proposition 5.9, we can find some $t_{k}\in\mathbb{R}$ such that $\nu_{t_{k}}^{k,\uparrow}\leq_{\mathrm{st}}\omega(B(\delta_{a_{k}},r_{k}))\leq_{\mathrm{st}}\nu_{t_{k}}^{k,\downarrow}$. Theorem 5.7 (iii) suggests that $\sup_{\mu\in B(\delta_{a_{k}},r_{k})}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$.
@@
(v). Noticing the conditions in Theorem 4.2 are verified in the proof of Theorem 2.1, we apply Theorem 4.2 (iv) to finish the proof. If $\mu\leq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{1}},r_{1})$, then from Theorem 4.2 (iv), we have $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$. If $\mu\geq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{n}},r_{n})$, the convergence result follows as well.
@@
On the other hand, Theorem 4.3 yields
@@
then Theorem 4.2 (v) shows that $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{1})\to 0$ as $t\to\infty$.
@@
Case $k=n$: Let $\{\nu^{n-1,\uparrow}_{t}\}_{t\in\mathbb{R}}$ be the increasing connecting orbit from $\nu_{2n-2}=\mu_{m_{2n-2}}$ to $\nu_{2n-1}$ and fix any $m\in(m_{2n-2},\infty)$. Similar to Case $k=1$, there exists $t_{0}\in\mathbb{R}$ such that $\mu_{m}\geq_{\mathrm{st}}\nu^{n-1,\uparrow}_{t_{0}}$. Then applying Theorem 4.2 (v) again, we conclude $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{2n-1})\to 0$ as $t\to\infty$. ∎
6 Applications to Classical Examples: Proof of Theorem 2.5-2.9
@@
According to Theorem 2.3, it suffices to show that equation (2.4) are locally dissipative at $-1,1$ with configurations $(r,g_{-1}),(r,g_{1})$ for $r=\frac{9-\sqrt{17}}{8}$ and for some $g_{\pm 1}$.
@@
$-2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big{(}2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}\big{)},$
@@
Then by calculation, we know that for any $w>0$, $g_{1}(\cdot,w)$ is convex if and only if
@@
By (6.2), we know that $g_{1}(\cdot,r^{2})$ is convex. Hence,
@@
Hence, for any $10<r<$ 1 and choosing $g_{1}$ as in (6.1), equation (2.4) is locally dissipative at $1$ with configuration $(r,g_{1})$ whenever
@@
Construction of $g_{-1}$: Similarly, let $g_{-1}=g_{1}$ be given as in (6.1), we can show that equation (2.4) is also locally dissipative at $-1$ with configuration $\big{(}\frac{9-\sqrt{17}}{8},g_{-1}\big{)}$. ∎
Proof of Theorem 2.6.
By a simple calculation, we know that
$-2xV^{\prime}(x+1)\leq-2x^{2}(x+1)(x+2)+\frac{1}{4}.$
Then we have
$-2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big{(}2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}-\frac{1}{4}\big{)}.$
Choose $g_{1}$ as follows: for any $z,w\geq 0$,
Repeating the proof of that in Theorem 2.5, we know that the equation (2.6) is locally dissipative at $1$ with configuration $(r,g_{1})$ with $r=\frac{9-\sqrt{17}}{8}$ whenever
i.e., (2.7) holds. Similarly, choose $g_{-1}=g_{1}$ as in (6.5), the equation (2.6) is locally dissipative at $-1$ with configuration $\big{(}\frac{9-\sqrt{17}}{8},g_{-1}\big{)}$. Then the desired results follow from Theorem 2.3 directly. ∎
@@
$g_{1}(z,w)$ $=2z^{2}-8z^{\frac{3}{2}}+(6+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{1}$ $=\frac{6-3\sqrt{3}}{2},$
$g_{-2}(z,w)$ $=2z^{2}-10z^{\frac{3}{2}}+(12+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{-2}$ $=\frac{(\sqrt{33}-15)^{2}}{64}.$
Through calculations similar in Theorem 2.5, local dissipative conditions hold at $1,-2$ for the equation (2.8). We apply Theorem 2.3 to get the desired results. ∎
@@
Take
$g_{0}(z,w)$ $=2z^{3}-10z^{2}+(8+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{0}$ $=\frac{5-\sqrt{13}}{3},$
$g_{2}(z,w)$ $=2z^{3}-20z^{\frac{5}{2}}+70z^{2}-100z^{\frac{3}{2}}+(48+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{2}$ $=r_{0},$
$g_{-2}(z,w)$ $=g_{2}(z,w),$ $r_{-2}$ $=r_{0}.$
Then the equation (2.9) is locally dissipative at $0,\pm 2$. The proof is finished by applying Theorem 2.3. ∎
Proof of Theorem 2.9.
The conclusions directly follow from Theorem 2.1. ∎

### v2

@@ 4.3 Proof of Main Theorems
(iv). For $k=1$, according to the statement (ii), the open ball $B(\delta_{a_{1}},r_{1})$ is positively invariant under $P^{*}$. By Theorem 1.4 (i)(ii) and Remark 4.1, there is only one invariant measure $\nu_{1}$ in $\overline{B(\delta_{a_{1}},r_{1})}$. Theorem 1.2 (ii) and Theorem 1.4 (i) entail that $\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{1}},r_{1})$ and hence $\omega(B(\delta_{a_{1}},r_{1}))\subset[\nu_{1},\nu_{2n-1}]$. Then $\nu_{1}\leq_{\mathrm{st}}\omega(B(\delta_{a_{1}},r_{1}))$. Let $\{\nu_{t}^{1,\downarrow}\}_{t\in\mathbb{R}}$ be a decreasing connecting orbit from $\nu_{2}$ to $\nu_{1}$ given in Theorem 1.4 (iii). Proposition 4.8 shows that there exists $t_{1}\in\mathbb{R}$ such that $\nu_{t_{1}}^{1,\downarrow}\geq_{\mathrm{st}}\omega(B(\delta_{a_{1}},r_{1}))$. Note that, by [36, Theorem 3.5], the semigroup $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is an order-preserving semiflow. Then Theorem 4.4 (ii) implies that $\nu_{1}$ attracts $B(\delta_{a_{1}},r_{1})$ under $P^{*}$, which means $\sup_{\mu\in B(\delta_{a_{k}},r_{k})}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$ as $t\to\infty$.
For $k=2,3,\dots,n-1$, let $\{\nu_{t}^{k,\downarrow}\}_{t\in\mathbb{R}}$ be a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$ and $\{\nu_{t}^{k,\uparrow}\}_{t\in\mathbb{R}}$ be an increasing connecting orbit from $\nu_{2k-2}$ to $\nu_{2k-1}$ given in Theorem 1.4 (iii). It follows from the statement (ii) that $B(\delta_{a_{k}},r_{k})$ is positively invariant under $P^{*}$. By Theorem 1.4 (i)(ii) and Remark 4.1, there is only one invariant measure $\nu_{2k-1}$ in $\overline{B(\delta_{a_{k}},r_{k})}$. We also see the order interval $\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{k}},r_{k})$ by Theorem 1.2 (ii) and Theorem 1.4 (i). By Proposition 4.8, we can find some $t_{k}\in\mathbb{R}$ such that $\nu_{t_{k}}^{k,\uparrow}\leq_{\mathrm{st}}\omega(B(\delta_{a_{k}},r_{k}))\leq_{\mathrm{st}}\nu_{t_{k}}^{k,\downarrow}$. Theorem 4.4 (iii) suggests that $\sup_{\mu\in B(\delta_{a_{k}},r_{k})}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$.
@@
(v). Noticing the conditions in Proposition 3.2 are verified in the proof of Theorem 1.2, we apply Proposition 3.2 (iv) to finish the proof. If $\mu\leq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{1}},r_{1})$, then from Proposition 3.2 (iv), we have $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$. If $\mu\geq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{n}},r_{n})$, the convergence result follows as well.
@@
On the other hand, Lemma 4.5 yields
@@
then Proposition 3.2 (v) shows that $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{1})\to 0$ as $t\to\infty$.
@@
Case $k=n$: Let $\{\nu^{n-1,\uparrow}_{t}\}_{t\in\mathbb{R}}$ be the increasing connecting orbit from $\nu_{2n-2}=\mu_{m_{2n-2}}$ to $\nu_{2n-1}$ and fix any $m\in(m_{2n-2},\infty)$. Similar to Case $k=1$, there exists $t_{0}\in\mathbb{R}$ such that $\mu_{m}\geq_{\mathrm{st}}\nu^{n-1,\uparrow}_{t_{0}}$. Then applying Proposition 3.2 (v) again, we conclude $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{2n-1})\to 0$ as $t\to\infty$. ∎
Before proving Theorem 1.1, we first give the following key proposition to verify the local dissipativity condition.
Proposition 4.9.
Assume that $W(x)=\frac{\theta}{2}x^{2}$ and the confining potential function $V\in C^{2}(\mathbb{R};\mathbb{R})$ satisfy the following conditions:
  1. (i)
    There exists $a\in\mathbb{R}$ such that $V^{\prime}(a)=0$ and $V^{\prime\prime}(a)>0$;
  2. (ii)
    $\liminf_{|x|\to\infty}\frac{V^{\prime}(x)}{x|x|^{\delta}}>0$ for some $\delta>0$.
Then, there exists $r_{a}>0$ such that for any $0<r\leq r_{a}$, there are $\theta_{r}>0$ and $\sigma_{r}>0$ such that for all $\theta\geq\theta_{r}$ and $\sigma^{2}(x)\leq\sigma_{r}$ for all $x\in\mathbb{R}$, the equation (1.1) is locally dissipative at $a$ with radius $r$.
Proof.
Without loss of generality, we assume $0<\delta<1$. Then we first show that there exist $C_{1},C_{2},C_{3}>0$ such that
$\frac{V^{\prime}(x+a)}{x}\geq C_{1}|x|^{\delta}-C_{2}|x|^{\frac{\delta}{2}}+C_{3},\ \ \text{for any $x\neq 0$}.$ (4.20)
By (ii), it holds that $\liminf_{|x|\to\infty}\frac{V^{\prime}(x+a)}{x|x|^{\delta}}>0$. Thus, there exists $R>0$ and $C_{1}>0$ such that
$\frac{V^{\prime}(x+a)}{x}\geq C_{1}|x|^{\delta},\ \text{ for all }\ |x|\geq R.$
Let $C_{3}=\frac{V^{\prime\prime}(a)}{2}>0$. Notice that $\lim_{x\to 0}\frac{V^{\prime}(x+a)}{x}=V^{\prime\prime}(a)>0$, there exist $0<\epsilon_{a}<R$ and such that
$\frac{V^{\prime}(x+a)}{x}\geq C_{3},\ \text{ for all }\ |x|\leq\epsilon_{a}.$
Take
$C_{2}=\epsilon_{a}^{-\frac{\delta}{2}}\bigg(\sup_{\epsilon_{a}\leq|x|\leq R}\bigg|\frac{V^{\prime}(x+a)}{x}\bigg|+C_{1}R^{\delta}+C_{3}\bigg).$
One can verify (4.20) in three cases: $\left\lvert x\right\rvert\geq R$, $\epsilon_{a}<|x|<R$ and $|x|\leq\epsilon_{a}$, separately. Then by (4.20) we have
$2xV^{\prime}(x+a)=-2|x|^{2}\frac{V^{\prime}(x+a)}{x}\leq-2C_{1}|x|^{2+\delta}+2C_{2}|x|^{2+\frac{\delta}{2}}-2C_{3}|x|^{2},\ \text{ for all }\ x\in\mathbb{R}.$ (4.21)
Let $\overline{\sigma}:=\sup_{x\in\mathbb{R}}|\sigma(x)|$. Choose
$g_{a}(z,w)=2C_{1}z^{1+\frac{\delta}{2}}-2C_{2}z^{1+\frac{\delta}{4}}+(2C_{3}+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2}.$
Obviously, $g_{a}$ is continuous and $g_{a}(z,\cdot)$ is decreasing for any $z\geq 0$ and hence
$\inf_{0\leq w\leq r^{2}}g_{a}(z,w)=g_{a}(z,r^{2}),\ \text{ for any }\ r>0.$ (4.22)
Also, (4.21) implies that
$2xV^{\prime}(x+a)-2x(W^{\prime}*\mu)(x)+|\sigma(x+a)|^{2}\leq-g_{a}(|x|^{2},\|\mu\|_{2}^{2}).$ (4.23)
For any $r>0$, a direct computation yields
$g_{a}(r^{2},r^{2})=2r^{2}\big(C_{1}r^{\delta}-C_{2}r^{\frac{\delta}{2}}+C_{3}\big)-\overline{\sigma}^{2},$
$\frac{\partial g_{a}(z,r^{2})}{\partial z}\bigg|_{z=r^{2}}=2\big(1+\frac{\delta}{2}\big)C_{1}r^{\delta}-2\big(1+\frac{\delta}{4}\big)C_{2}r^{\frac{\delta}{2}}+2C_{3}+\theta,$
and for all $z\geq 0$,
$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}=\delta\big(1+\frac{\delta}{2}\big)C_{1}z^{-(1-\frac{\delta}{2})}-\frac{\delta}{2}\big(1+\frac{\delta}{4}\big)C_{2}z^{-(1-\frac{\delta}{4})}+\frac{\theta r}{2}z^{-\frac{3}{2}}.$ (4.24)
Then there exists $r_{a}>0$ such that for any $0<r\leq r_{a}$, we have
$2r^{2}\big(C_{1}r^{\delta}-C_{2}r^{\frac{\delta}{2}}+C_{3}\big)>0,$ (4.25)
and
$\frac{\partial g_{a}(z,r^{2})}{\partial z}\bigg|_{z=r^{2}}=2\big(1+\frac{\delta}{2}\big)C_{1}r^{\delta}-2\big(1+\frac{\delta}{4}\big)C_{2}r^{\frac{\delta}{2}}+2C_{3}+\theta>0.$ (4.26)
Now, we fix $0<r\leq r_{a}$. By (4.24), we have $\lim_{z\downarrow 0}\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}=\infty$. Hence, there exists $\epsilon_{r}>0$ such that
$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}>0,\ \text{ for all }\ 0<z<\epsilon_{r}.$
On the other hand,
$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}=\frac{\delta}{8}z^{(\frac{\delta}{4}-1)}\big(4(2+\delta)C_{1}z^{\frac{\delta}{4}}-(4+\delta)C_{2}\big)+\frac{\theta r}{2}z^{-\frac{3}{2}}>0,\ \text{ for all }\ z\geq C_{\delta}:=\Big(\frac{(4+\delta)C_{2}}{4(2+\delta)C_{1}}\Big)^{\frac{4}{\delta}}.$
Let
$\theta_{r}=\frac{\delta}{r}\big(1+\frac{\delta}{4}\big)C_{2}C_{\delta}^{\frac{3}{2}}\epsilon_{r}^{-(1-\frac{\delta}{4})}.$ (4.27)
For $\theta\geq\theta_{r}$, by (4.24), one has
$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}\geq\delta\big(1+\frac{\delta}{2}\big)C_{1}z^{-(1-\frac{\delta}{2})}>0,\ \text{ for all }\ \epsilon_{r}\leq z\leq C_{\delta}.$
Thus, for all $\theta\geq\theta_{r}$, we conclude that
$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}>0,\ \text{ for all }\ z>0.$ (4.28)
So $g_{a}(\cdot,r^{2})$ is convex. Thus, by (4.26),
$\frac{\partial g_{a}(z,r^{2})}{\partial z}\geq\frac{\partial g_{a}(r^{2},r^{2})}{\partial z}>0,\ \text{ for all }\ z\geq r^{2}.$ (4.29)
(4.25) implies that there exists $\sigma_{r}$ such that
$0<\sigma_{r}<2r^{2}\big(C_{1}r^{\delta}-C_{2}r^{\frac{\delta}{2}}+C_{3}\big).$ (4.30)
Then by (4.29), we have
$g_{a}(z,r^{2})\geq g_{a}(r^{2},r^{2})>\sigma_{r}-\overline{\sigma}^{2}\geq 0,\ \text{ for all }\ z\geq r^{2}.$ (4.31)
Therefore, for any $0<r\leq r_{a}$, let $\theta_{r},\sigma_{r}$ be given as in (4.27) and (4.30). It follows from (4.22), (4.23), (4.28), (4.31) that for all $\theta\geq\theta_{r},\ \overline{\sigma}^{2}\leq\sigma_{r}$, equation (1.1) is locally dissipative at $a$ with radius $r$. ∎
Remark 4.2.
If $V^{\prime}$ is hyper-dissipative (see Assumption 1 (iii)), i.e., there exist $\alpha>0$, $\beta>0$, $\delta>0$ such that, for all $x,y\in\mathbb{R}$,
$(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq-\alpha\left\lvert x-y\right\rvert^{2+\delta}+\beta.$
Then we have
$xV^{\prime}(x)\leq-\alpha\left\lvert x\right\rvert^{2+\delta}+|x||V^{\prime}(0)|+\beta,$
and hence
$\frac{V^{\prime}(x)}{x|x|^{\delta}}=\frac{xV^{\prime}(x)}{|x|^{2+\delta}}\geq\alpha-\frac{|V^{\prime}(0)|}{|x|^{1+\delta}}-\frac{\beta}{|x|^{2+\delta}}.$
Thus, $\lim_{|x|\to\infty}\frac{V^{\prime}(x)}{x|x|^{\delta}}=\alpha>0$ and the condition (ii) in Proposition 4.9 holds true.
@@
According to Remark 4.2, we have
$\lim_{x\to\infty}V^{\prime}(x)=\infty,\ \ \lim_{x\to-\infty}V^{\prime}(x)=-\infty.$ (4.32)
Since $V^{\prime}$ has only finitely many zeros and each of them is simple, the number of zeros must be odd. Denoting them by $a_{1}<b_{1}<a_{2}<b_{2}<\cdots<a_{n-1}<b_{n-1}<a_{n}$, it then follows that
$V^{\prime\prime}(a_{i})>0,\ 1\leq i\leq n,\ \text{ and }\ V^{\prime\prime}(b_{j})<0,\ 1\leq j\leq n-1.$ (4.33)
By virtue of Proposition 4.9, Remark 4.2 and (4.33) that, there exist
$0<r<\frac{1}{2}\min_{1\leq i\leq n-1}\left\lvert a_{i+1}-a_{i}\right\rvert,\ \ \theta_{0}>0,\ \text{ and }\ \sigma_{0}>0$
such that, for all $\theta\geq\theta_{0}$ and $\overline{\sigma}\leq\sigma_{0}$, equation (1.1) is locally dissipative at $a_{k}$ with radius $r$ for all $1\leq k\leq n$. Then the desired result follows from Theorem 1.4. ∎
Proof of Theorem 1.6.
According to Theorem 1.4, it suffices to show that equation (1.9) are locally dissipative at 1,1-1,1 with configurations $(r,g_{-1}),(r,g_{1})$ for $r=\frac{9-\sqrt{17}}{8}$ and for some $g_{\pm 1}$.
@@
$2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big(2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}\big),$
@@
Then by calculation, we have for any $w>0$, $g_{1}(\cdot,w)$ is convex if and only if
@@
By (4.35), we have $g_{1}(\cdot,r^{2})$ is convex. Hence,
@@
Hence, for any $0<r<1$ and choosing $g_{1}$ as in (4.34), equation (1.9) is locally dissipative at $1$ with configuration $(r,g_{1})$ whenever
@@
Construction of $g_{-1}$: Similarly, let $g_{-1}=g_{1}$ be given as in (4.34), we can show that equation (1.9) is also locally dissipative at 1-1 with configuration $\big(\frac{9-\sqrt{17}}{8},g_{-1}\big)$. ∎
@@
$g_{0}(z,w)$ $=2z^{3}-10z^{2}+(8+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{0}$ $=\frac{\sqrt{15-3\sqrt{13}}}{3},$
$g_{2}(z,w)$ $=2z^{3}-20z^{\frac{5}{2}}+70z^{2}-100z^{\frac{3}{2}}+(48+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{2}$ $=r_{0},$
$g_{-2}(z,w)$ $=g_{2}(z,w),$ $r_{-2}$ $=r_{0},$
where $\overline{\sigma}:=\sup_{x\in\mathbb{R}}|\sigma(x)|$. Similar to the calculation in the proof of Theorem 1.6, we can prove that the equation (1.11) is locally dissipative at $0,\pm 2$. The proof is finished by applying Theorem 1.4. ∎
@@
By a simple calculation, we have
$2xV^{\prime}(x+1)\leq-2x^{2}(x+1)(x+2)+\frac{1}{4}.$
Then we have
$2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big(2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}-\frac{1}{4}\big),$
where $\overline{\sigma}:=\sup_{x\in\mathbb{R}}|\sigma(x)|$. Choose $g_{1}$ as follows: for any $z,w\geq 0$,
Repeating the proof of that in Theorem 1.6, we have the equation (1.12) is locally dissipative at $1$ with configuration $(r,g_{1})$ with $r=\frac{9-\sqrt{17}}{8}$ whenever
i.e., (1.13) holds. Similarly, choose $g_{-1}=g_{1}$ as in (4.38), the equation (1.12) is locally dissipative at 1-1 with configuration $\big(\frac{9-\sqrt{17}}{8},g_{-1}\big)$. Then the desired results follow from Theorem 1.4 directly. ∎
Motivated by our results, we conclude the paper with the following conjecture.
In this work, by exploiting the order-preserving property of the associated semiflow, we have obtained a partial characterization of the basins of attraction of invariant measures, both in the sense of the stochastic order and the topology. In particular, our results reveal the existence of unbounded open subsets contained in the basins of attraction of minimal/maximum invariant measures.
In the theory of monotone dynamical systems, a classical saddle-point structure (or saddle-type behavior) arises when a system possesses three equilibria $a<b<c$, where $a$ and $b$ are stable. More precisely, the state space can be decomposed into three disjoint invariant sets: the basin of attraction $B_{1}$ of $a$, the basin of attraction $B_{2}$ of $b$, and an unordered invariant set $M$ containing $c$. The set $M$ separates the two basins $B_{1}$ and $B_{2}$ and is usually referred to as the separatrix (see e.g., Jiang-Liang-Zhao [31]).
Conjecture.
In the double-well granular media equation, the three invariant measures admit an analogous saddle-point structure. More precisely, the unstable invariant measure $\mu_{0}$ is expected to be contained in an unordered invariant separating set that acts as a separatrix between the basins of attraction of the two stable invariant measures.